OpenStudy (kobeni-chan):

Fan and Medal! :) Sandra is riding the Ferris wheel, and her height can be modeled by the equation H(t) = 25 cos[(pi/14)t] + 31, where H represents the height of the person above the ground in feet at t seconds. How far above the ground is Sandra before the ride begins?

OpenStudy (kobeni-chan):

I thought that I would be substituting t for 0, and then solving, but I got stuck >_< Right now I have 25 cos(0) + 31

OpenStudy (kobeni-chan):

@ganeshie8 @Tootles143 @pooja195 Please can someone help me? :(

OpenStudy (kobeni-chan):

@77777jeannie77777 @cupcakerain @Gokuporter @Awolflover1

OpenStudy (kobeni-chan):

@amilapsn @Omar_Elboredy

OpenStudy (anonymous):

cos(0) = 1 so assign it and you have the height

OpenStudy (anonymous):

So .. What before the ride begins mean ? It means that just starting the ride which implies that it is wanted to get the height at time(t) = 0 then substituting back into the equation with this time you will get : H(0) = 25 cos( (π/14)*0 ) + 31 = 25 cos(0) + 32 = ? Solve it Urself ! :D Let me know if you got it :)

OpenStudy (anonymous):

*31

OpenStudy (kobeni-chan):

I got 25 32. But I don't know what to do about the space :/

OpenStudy (anonymous):

25+ 32

OpenStudy (anonymous):

Your equation after subs. by t = 0 will be sth like that : 25 cos(0) + 31 = .... Ok ?

OpenStudy (anonymous):

It's 31 not 32 , I re-corrected it ! :D

OpenStudy (kobeni-chan):

Wait so do I add them to get the height or do I leave it as an equation??

OpenStudy (anonymous):

You have to add them to get the height

OpenStudy (anonymous):

yeah t=0 H(0) = 25+32 = 57 which is the height at t=0

OpenStudy (kobeni-chan):

Ok thank you guys :) And since it doesn't give a unit of measure, I just say 57 units, right?

OpenStudy (anonymous):

"H represents the height of the person above the ground in ""feet"" " The given unit is "feet" ! :) So the answer will be : 57 feet above the ground

OpenStudy (kobeni-chan):

Oh I can't believe I missed that >_< thanks again!